OpenStudy (anonymous):

Can someone please show me how to solve this problem step-by-step? x^2-2x+8=0

OpenStudy (anonymous):

\[a. 1\pm i \sqrt{7}\] \[b. \pm i \sqrt{7}\] \[c. 1\pm 2i \sqrt{7}\]

OpenStudy (anonymous):

\[x= \frac{- ( 2 ) \pm \sqrt{ (-2)^2 - ( 4 )( 1 ) ( 8 )} }{ 2 ( 1 )} \]

OpenStudy (anonymous):

ohhh you use the quadratic equation!